# Bandit Sequential Posted Pricing via Half-Concavity

### v1

@@ 3.3.2 Step 1: Finding $(\hat{p}_{1},\cdots,\hat{p}_{n})$ to Approximate $(p^{*}_{1},\cdots,p^{*}_{n})$
The proof of 3.6 is almost identical to 2.4. We leave the details to Section B.1. The claim shows that $R_{i}(p)$ is half-concave, and therefore learnable in $[\mathsf{Rev}(S_{i+1}),1]$. However, in $[0,\mathsf{Rev}(S_{i+1})]$ the function is increasing but not necessarily concave. The subroutine $\mathsf{SubR}$ will involve different cases to handle non-concavity inside interval $[0,\mathsf{Rev}(S_{i+1})]$. Depending on the value of $P_{i}$ and $F_{i}(r^{s}_{i})$, we run different algorithms. See Algorithm 5 for the pseudo-code of $\mathsf{SubR}$.
@@
5Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=r^{s}_{i}$ to estimate $\hat{F}_{i}(r^{s}_{i}):=\frac{\sum\mathbf{1}[\text{Item not sold before }i+1]}{\sum\mathbf{1}[\text{Item not sold before }i]}$.
@@
ALGORITHM 5 Subroutine $\mathsf{SubR}$
@@
Case 1 ($P_{i}$ is small): Since the difference between $R_{i}(\hat{p}^{*}_{i})$ and any other $R_{i}(p)$ can be bounded by $P_{i}$, any price in $[\ell_{i},r^{s}_{i}]$ is feasible to become $\hat{p}_{i}$.
Case 2 ($F_{i}(r^{s}_{i})$ is large): In this case, we will show that the value of $\mathsf{Rev}(S_{i+1})$. can be estimated accurately. Then, we divide interval $[\ell_{i},r^{s}_{i}]$ from the upper bound of $\mathsf{Rev}(S_{i+1})$ into two parts. For the part on the right, 3.6 guarantees that $R_{i}(p)$ is half-concave in this part, and we call the single-buyer algorithm to find a candidate of $\hat{p}_{i}$. For the part on the left, the constrained Lipschitzness from 3.6 suggests that we can mark the lower bound of $\mathsf{Rev}(S_{i+1})$ as a candidate price. Finally, we take a better one to be $\hat{p}_{i}$ from two candidate prices.
@@
3.7 suggests that $R_{i}(p)$ is learnable when $F_{i}(r^{s}_{i})$ is small and $r^{s}_{i}\geq\mathsf{Rev}(S_{i+1})$ holds. The proof of 3.7 requires ideas of generalized half-concavity. We defer the details to Section 3.3.4.
@@
Suppose $\hat{p}^{*}_{i}<\mathsf{Rev}(S_{i+1})$. Then, 3.5 guarantees that $p^{*}_{i}\geq\mathsf{Opt}(S_{i+1})\geq\mathsf{Rev}(S_{i+1})>\hat{p}^{*}_{i}$. Since we have $\hat{p}^{*}_{i},p^{*}_{i}\in[\ell_{i},r_{i}]$, there must be $\mathsf{Rev}(S_{i+1})\in[\ell_{i},r_{i}]$. This leads to a contradiction: 3.6 guarantees $R_{i}(p)$ is increasing in $[0,\mathsf{Rev}(S_{i+1})]$, which is in contrast to the assumption that $\hat{p}^{*}_{i}<\mathsf{Rev}(S_{i+1})$. Therefore, there must be $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})$. ∎
@@
Now, we give the formal proof of 3.4.
Proof of 3.4.
The bound of number of tested rounds in and the constraint that $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$ are directly guaranteed by Algorithm 5. It remains to show $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. We start the proof by first providing the following concentration bounds for $\mathsf{SubR}$:
@@
Next, we prove $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ separately for the following cases:
@@
Case 2: $\hat{P}_{i}\geq\frac{3}{4}\delta$ while $\hat{F}_{i}(r^{s}_{i})\geq 0.4$ (Line 7-12). Consider the value of $\hat{p}^{*}_{i}$. If $\hat{p}^{*}_{i}\in[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$, there must be $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})\leq\delta$. This is true because 3.6 guarantees $[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]\subseteq[\mathsf{Rev}(S_{i+1}),1]$ is half-concave, and 2.7 guarantees that the single-buyer algorithm outputs $\hat{p}^{(1)}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})\leq\delta$ with probability $1-T^{-6}$.
If $\hat{p}^{*}_{i}\notin[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$, there must be $\hat{p}^{*}_{i}\in[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})-\frac{\delta}{\hat{P}_{i}}\},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\}]$. This is true because simultaneously we have $\hat{p}^{*}_{i}\geq\ell_{i}$ from definition of $\hat{p}^{*}_{i}$, $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})\geq\widehat{\mathsf{Rev}}(S_{i+1})-\frac{\delta}{\hat{P}_{i}}$ from 3.8 and 3.9, and $\hat{p}^{*}_{i}\leq\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}$ from the assumption $\hat{p}^{*}_{i}\notin[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$. Then, the constrained Lipschitzness from 3.6 gives
@@
In this case, the algorithm prepares two candidate prices. If $\mathsf{Rev}(S_{i+1})\leq r^{s}_{i}$, 3.7 guarantees that with probability $1-T^{-6}$ we get $\hat{p}^{(3)}_{i}$ such that $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(3)}_{i})\leq\delta$. Otherwise, there must be $\mathsf{Rev}(S_{i+1})>r^{s}_{i}$. In this case, 3.8 guarantees $r^{s}_{i}<\mathsf{Rev}(S_{i+1})\leq\hat{p}^{*}_{i}$, and the constrained Lipschitzness of $R_{i}(p)$ guarantees
@@
To conclude, $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ holds in all cases. The final step is to check the success probability of our proof. In this proof, we require the correctness of 3.9, Algorithm 2 in Case 2, and $\mathsf{Alg}$ in Case 3. Each of them fails with probability $T^{-6}$. By the union bound, the whole proof succeeds with probability $1-3T^{-6}>1-T^{-5}$. ∎
3.3.4 3.7 via Generalized Half-Concavity
In this subsection, we prove the missing 3.7.
See 3.7
Our algorithm begins from the following observation: Notice that
@@
In fact, function $R_{i}(p)$ is half 2-concave in any interval $[\ell,r]\subseteq[0,1]$ such that $[\ell,r]\cap[\mathsf{Rev}(S_{i+1}),1]\neq\emptyset$. Here we only prove the half 2-concavity on $[\ell_{i},r_{i}]$ to satisfy the need of 3.7.
@@
(i) Single-peakedness: This is guaranteed by 3.6.
(ii) Lipschitzness with respect to $p^{*}$: 3.8 gives $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})$. Then, the constrained Lipschitzness from 3.6 is sufficient to give the inequality $R_{i}(\hat{p}^{*}_{i})-R_{i}(p)\leq\hat{p}^{*}_{i}-p$ for all $p\in[\ell_{i},\hat{p}^{*}_{i}]$.
@@
  • Fact 2: For $x\leq\mathsf{Rev}(S_{i+1})$ and $y>\mathsf{Rev}(S_{i+1})$, the constrained Lipschitzness in 3.6 gives
@@
3.12 is a generalization of Algorithm 2 for a single buyer. Observe that Algorithm 2 only needs a weaker inequality $\frac{R(b)-R(a)}{b-a}\geq\frac{R(p^{*})-R(b)}{p^{*}-b}$. Then, generalizing Algorithm 2 by taking $\lambda$ into consideration gives the desired algorithm for 3.12. We leave the details of this generalization to Section B.3.
Finally, combining 3.12 with 3.11 directly proves 3.7.
3.3.5 Step 2: Determining New Confidence Intervals via Binary Search
@@
With 3.13, the idea of our binary search algorithm is clear: We first estimate $R_{i}(\hat{p}_{i})$ to be the benchmark. For a price $p$, we can estimate $R_{i}(p)$ and compare $R_{i}(p)+5(n-i)\delta$ with $R_{i}(\hat{p}_{i})$. If $R_{i}(p)+5(n-i)\delta<R_{i}(\hat{p}_{i})$, it is evident to say $R^{*}_{i}(p)<R^{*}_{i}(\hat{p}_{i})$, and therefore $p$ can’t be $p^{*}_{i}$. Otherwise, we keep $p$ as a candidate of $p^{*}_{i}$. Since function $R_{i}(p)$ is single-peaked, the experiment above can be done via two binary search algorithms, and extra steps are needed to take care of the errors from the uncertainty - we only have estimates $\hat{R}_{i}(p)$ and $\hat{R}_{i}(\hat{p}_{i})$, instead of $R_{i}(p)$ and $R_{i}(\hat{p}_{i})$. Here we only give the main lemma of our binary search algorithm and leave the details to Section B.4.
Lemma 3.14.
Assume 3.3 holds. Given hidden function $R_{i}(p)$, interval $[\ell_{i},r_{i}]\ni p^{*}_{i}$, error parameter $\delta>1/\sqrt{T}$, price $\hat{p}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$, there exists an algorithm that runs $O(\frac{\log^{2}T}{\delta^{2}})$ rounds with price vectors $(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}\in[\ell_{i},r^{s}_{i}],\hat{p}_{i+1},\cdots,\hat{p}_{n})$ and outputs with probability $1-T^{-5}$ $[\ell^{\prime}_{i},r^{\prime}_{i}]\subseteq[\ell_{i},r_{i}]$ that satisfies:
  • $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$.
Simply using 3.14 $n$ times to get $n$ new confidence intervals gives the second-step of our algorithm.
3.3.6 Finishing the Proof of 3.2
In 3.2, the bound for the number of rounds and the constraint that $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$ are guaranteed by 3.3 and 3.14. For the statement $p^{*}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$, 3.3 fails with probability $T^{-4}$. Each call of 3.14 fails with probability $T^{-5}$. By the union bound, with probability $1-T^{-4}-n\cdot T^{-5}>1-T^{-3}$ we get $[\ell^{\prime}_{1},r^{\prime}_{1}],\cdots,[\ell^{\prime}_{n},r^{\prime}_{n}]$ such that $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$ for all $i\in[n]$. To prove 3.2, it only remains to show every price vector $(p_{1},\cdots,p_{n})$ satisfying $p_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}-{T^{-100}}]$ has $\epsilon$ regret. We instead prove the following lemma to finish the proof of 3.2:
Lemma 3.15.
Assume 3.3 and 3.14 hold. For $p_{i},\cdots,p_{n}$ that satisfies $p_{j}\in[\ell^{\prime}_{j},r^{\prime}_{j}-{T^{-100}}]$ for all $j\in\{i,\cdots,n\}$, we have
Proof.
We prove the lemma via induction. The base case is $i=n$. 3.14 gives
For the induction step, we hope to show
under the induction hypothesis
Now we finish the induction step:
[…]
$+10(n-i+1)\delta+14\delta+50(n-i)^{2}\delta$
[…]
Here, the first inequality is from the inequality $P_{i}\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$ in 3.3. The second inequality is from 3.4 that guarantees $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. The third inequality is from 3.14 that guarantees $R_{i}(\hat{p}^{*}_{i})-R_{i}(p_{i})\leq 5(n-i)\delta+15\delta$. The fourth inequality is from the fact that prices $p^{*}_{i+1},\cdots,p^{*}_{n}$ are optimal for suffix $S_{i+1}$. The fifth inequality is from the corollary of the induction hypothesis. Comparing the start and the end of the above inequality finishes the induction step. ∎
Finally, taking $i=1$ in 3.15 with the fact that $\delta=\frac{\epsilon}{100n^{2}}$ finishes the proof of 3.2.

### v2

@@ 12       Let $r^{\prime}_{i}\leftarrow r_{i}$
Proof.
[…]
The above inequality shows $R_{i}(\mathsf{Rev}(S_{i+1}))$ is not worse than any price in $[0,\mathsf{Rev}(S_{i+1})]$. Therefore, there must be $\arg\max_{p}R_{i}(p)\geq\mathsf{Rev}(S_{i+1})$.
$\frac{R(c)}{P_{i}}~{}>~{}$ $\frac{R(b)}{P_{i}}+\frac{t\cdot(R(b)-R(a))}{P_{i}}$
[…]
$-t(t+1)(b-a)(q(a)-q(b))$
[…]
The claim shows that $R_{i}(p)$ is half-concave, and therefore learnable in $[\mathsf{Rev}(S_{i+1}),1]$. However, in $[0,\mathsf{Rev}(S_{i+1})]$ the function is increasing but not necessarily concave. The sub-algorithm $\mathsf{SubAlg}$ will involve different cases to handle non-concavity inside interval $[0,\mathsf{Rev}(S_{i+1})]$. Depending on the value of $P_{i}$ and $F_{i}(r^{s}_{i})$, we run different algorithms. See Algorithm 7 for the pseudo-code of $\mathsf{SubAlg}$.
@@
5
6Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=r^{s}_{i}$ to estimate $\hat{F}_{i}(r^{s}_{i}):=\frac{\sum\mathbf{1}[\text{Item not sold before }i+1]}{\sum\mathbf{1}[\text{Item not sold before }i]}$.
@@
ALGORITHM 7 Sub-algorithm $\mathsf{SubAlg}$
@@
Case 1 ($P_{i}$ is small): Since the difference between $R_{i}(\hat{p}^{*}_{i})$ and any other $R_{i}(p)$ is upper bounded by $P_{i}$, playing any price in $[\ell_{i},r^{s}_{i}]$ does not incur a big regret. Therefore, taking $\hat{p}_{i}$ to be an arbitrary price in $[\ell_{i},r^{s}_{i}]$ is feasible.
Case 2 ($F_{i}(r^{s}_{i})$ is large): In this case, the value of $\mathsf{Rev}(S_{i+1})$ can be estimated accurately. Then, we divide interval $[\ell_{i},r^{s}_{i}]$ from the upper bound of $\mathsf{Rev}(S_{i+1})$ into two parts. For the part on the right, 3.9 guarantees that $R_{i}(p)$ is half-concave, and we call the single-buyer algorithm to find a candidate price. For the part on the left, the constrained Lipschitzness from 3.9 suggests that we can mark the lower bound of $\mathsf{Rev}(S_{i+1})$ as a candidate price. Finally, we take a better one from two candidate prices to be $\hat{p}_{i}$.
@@
3.10 suggests that $R_{i}(p)$ is learnable when $F_{i}(r^{s}_{i})$ is small and $r^{s}_{i}\geq\mathsf{Rev}(S_{i+1})$ holds. The proof of 3.10 requires ideas of generalized half-concavity. For consistency, we first assume the correctness of 3.10, and defer the details to Section 3.8.
@@
Suppose $\hat{p}^{*}_{i}<\mathsf{Rev}(S_{i+1})$. Then, we have $p^{*}_{i}\geq\mathsf{Opt}(S_{i+1})\geq\mathsf{Rev}(S_{i+1})>\hat{p}^{*}_{i}$, where the first inequality is implied by the observation that
$R^{*}_{i}(p^{*}_{i})-R^{*}_{i}(\mathsf{Opt}(S_{i+1}))~{}=~{}P_{i}\cdot(1-F_{i}(p^{*}_{i}))\cdot(p^{*}_{i}-\mathsf{Opt}(S_{i+1}))~{}\geq~{}0.$
Since we have $\hat{p}^{*}_{i},p^{*}_{i}\in[\ell_{i},r_{i}]$, there must be $\mathsf{Rev}(S_{i+1})\in[\ell_{i},r_{i}]$. This leads to a contradiction: 3.9 guarantees $R_{i}(p)$ is increasing in $[0,\mathsf{Rev}(S_{i+1})]$, which is in contrast to the assumption that $\hat{p}^{*}_{i}<\mathsf{Rev}(S_{i+1})$. Therefore, there must be $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})$. ∎
@@
Now, we prove 3.5. Our proof starts from the following concentration bounds for $\mathsf{SubAlg}$:
@@
Proof.
where the last inequality holds when $C$ is a constant greater than $60$.
[…]
where the last inequality holds when $C$ is a constant greater than $16$. ∎
With the help of 3.12, we can give the formal proof of 3.5:
Proof of 3.5.
The bound of number of tested rounds in and the constraint that $p_{i}\in[\ell_{i},r^{s}_{i}]$ are directly guaranteed by Algorithm 7. It remains to show $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. We prove the inequality separately for the following cases:
@@
Case 2: $\hat{P}_{i}\geq\frac{3}{4}\delta$ while $\hat{F}_{i}(r^{s}_{i})\geq 0.4$ (Line 7-12). Consider the value of $\hat{p}^{*}_{i}$. If $\hat{p}^{*}_{i}\in[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$, there must be $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})\leq\delta$. This is true because 3.9 guarantees $[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]\subseteq[\mathsf{Rev}(S_{i+1}),1]$ is half-concave, and 2.7 guarantees that the single-buyer algorithm outputs $\hat{p}^{(1)}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})\leq\delta$ with probability $1-T^{-6}$.
If $\hat{p}^{*}_{i}\notin[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$, there must be $\hat{p}^{*}_{i}\in[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})-\frac{\delta}{\hat{P}_{i}}\},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\}]$. This is true because simultaneously we have $\hat{p}^{*}_{i}\geq\ell_{i}$ from definition of $\hat{p}^{*}_{i}$, $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})\geq\widehat{\mathsf{Rev}}(S_{i+1})-\frac{\delta}{\hat{P}_{i}}$ from 3.11 and 3.12, and $\hat{p}^{*}_{i}\leq\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}$ from the assumption $\hat{p}^{*}_{i}\notin[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$. Then, the constrained Lipschitzness from 3.9 gives
@@
In this case, the algorithm prepares two candidate prices. If $\mathsf{Rev}(S_{i+1})\leq r^{s}_{i}$, 3.10 guarantees that with probability $1-T^{-6}$ we get $\hat{p}^{(3)}_{i}$ such that $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(3)}_{i})\leq\delta$. Otherwise, there must be $\mathsf{Rev}(S_{i+1})>r^{s}_{i}$. In this case, 3.11 guarantees $r^{s}_{i}<\mathsf{Rev}(S_{i+1})\leq\hat{p}^{*}_{i}$, and the constrained Lipschitzness of $R_{i}(p)$ guarantees
@@
To conclude, $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ holds in all cases. The final step is to check the success probability of our proof. In this proof, we require the correctness of 3.12, Algorithm 2 in Case 2, and $\mathsf{Alg}$ in Case 3. Each of them fails with probability $T^{-6}$. By the union bound, the whole proof succeeds with probability $1-3T^{-6}>1-T^{-5}$. ∎
3.8 3.10 via Generalized Half-Concavity
In this subsection, we prove the missing 3.10, which is restated for convenience.
See 3.10
Our algorithm begins from the following observation: Recall that $R_{i}(p)=\mathsf{Pre}(S_{i})+P_{i}(p\cdot(1-F_{i}(p))+\mathsf{Rev}(S_{i+1})\cdot F_{i}(p))$. Then,
@@
The intuition of defining half $\lambda$-concavity is to capture the case when $R(p)$ is concave, or the case when $R(p)$ has slope between $[k,\lambda k]$, or a mixture of these two. The following lemma suggests $R_{i}(p)$ is captured by this half $\lambda$-concavity when $F_{i}(r^{s}_{i})$ is not too large.
@@
(i) Single-peakedness: This is guaranteed by 3.9.
(ii) Lipschitzness with respect to $p^{*}$: 3.11 gives $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})$. Then, the constrained Lipschitzness from 3.9 is sufficient to give the inequality $R_{i}(\hat{p}^{*}_{i})-R_{i}(p)\leq\hat{p}^{*}_{i}-p$ for all $p\in[\ell_{i},\hat{p}^{*}_{i}]$.
@@
  • Fact 2: For $x\leq\mathsf{Rev}(S_{i+1})$ and $y>\mathsf{Rev}(S_{i+1})$, the constrained Lipschitzness in 3.9 gives
@@
3.15 is a generalization of Algorithm 2 for a single buyer. Observe that Algorithm 2 only needs a weaker inequality $\frac{R(b)-R(a)}{b-a}\geq\frac{R(p^{*})-R(b)}{p^{*}-b}$. Then, generalizing Algorithm 2 by taking $\lambda$ into consideration gives the desired algorithm for 3.15.
Observe that combining 3.15 with 3.14 directly proves 3.10. It only remains to prove 3.15.
1 Fix $p_{j}=r^{s}_{j}$ for $j<i$ and $p_{j}=\hat{p}_{j}$ for $j>i$.
3 while $r_{j}-\ell_{j}>\delta$ do
7       if $\hat{R}_{i}(a_{j})<\hat{R}_{i}(b_{j})-2\delta$ then
8            Let $\ell_{j+1}\leftarrow a_{j},r_{j+1}\leftarrow r_{j}$
9      else
10            Let $\ell_{j+1}\leftarrow\ell_{j},r_{j+1}\leftarrow b_{j}$
11      $j\leftarrow j+1$
Output: $\hat{p}_{i}$.
Proof of 3.15.
Algorithm 8 is the desired algorithm for 3.15. The algorithm is almost the same as Algorithm 2. The only difference is that extra care is needed for the extra parameter $\lambda$.
@@ 11      $j\leftarrow j+1$
Now, we prove $R_{i}(p^{*}_{\ell,r})-R_{i}(\hat{p}_{i})\leq 6\lambda\delta$ assuming 3.16 holds. We demonstrate the proof by discussing the following two cases. The first case is that there exists $j\in[k]$, such that $p^{*}_{\ell,r}$ falls in $[\ell_{j},r_{j}]$ but not in $[\ell_{j+1},r_{j+1}]$. In this case, there must be $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ and $[b_{j},r_{j}]$ is dropped. If not, we have $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ and $[\ell_{j},a_{j}]$ is dropped. However, the inequality $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ together with 3.16 gives $R(a_{j})<R(b_{j})$. Then, the single-peakedness of $R_{i}(p)$ gives $p^{*}_{\ell,r}\geq a_{j}$, which is in contrast to the assumption that $p^{*}_{\ell,r}\notin[\ell_{j+1},r_{j+1}]$. Therefore, we have $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$, and the assumption $p^{*}_{\ell,r}\notin[\ell_{j+1},r_{j+1}]$ implies $p^{*}_{\ell,r}\in[b_{j},r_{j}]$. Then, the half $\lambda$-concavity of $R_{i}(p)$ in $[a_{j},p^{*}_{\ell,r}]$ guarantees